# Sloth: scaling laws for LLM skills to predict multi-benchmark performance across families

### v4

@@ 5 Conclusion
We include a paragraph about limitations in Appendix A.
Appendix A Limitations
From the predictive side of Sloth, we believe that the main limitation is that the model is still dependent, most of the time, on seeing data from at least one LLM from the LLM family of interest. Moreover, we train the link function in the best version of Sloth using flexible neural networks, which can interpolate data very well, but have no guarantee of extrapolation when the (hypothetical) LLM of interest is very different from others in the training set. From the interpretability side, we only understand the identifiability problems, such as transformations in the latent space, that can arise in the most simple case of Sloth. This fact limits our understanding and interpretability of the most advanced versions of the model.
Appendix B Identifiability of model parameters
To interpret Sloth parameters, we need to guarantee they are identifiable. Given that our scaling law models the function $\mu_{ij}(s,t)={\mathbf{E}}[Y_{ij}(s,t)]$, that condition is equivalent to the following statement: if two sets of parameters are responsible for characterizing $\mu_{ij}(s,t)$, then those set of parameters should be the same up to predictable variations such as translations or rotations. To prove identifiability, we work with a fixed and invertible $\sigma$, as usually done in the literature, and assume $\gamma_{j}$’s are fixed. The last condition is reasonable since these constants are usually known beforehand, e.g., it is well accepted that the lower asymptote $\gamma_{j}$ for MMLU (Hendrycks et al., 2020) performance is $25\%$ which is given by $100\%$ divided by the number of multiple-choice alternatives. Denote our fixed design matrix as $X\in{\mathbf{R}}^{n\times p}$, where each row is given by an LLM and $p$ equals 3 plus the number of families, and define
@@ Appendix B Identifiability of model parameters
such that the rows of $XB\in{\mathbf{R}}^{n\times d}$ give the skills vectors $\theta^{(i)}\triangleq(XB)^{(i)}$’s of all models in our dataset. Here $n$ denotes the total number of models in the dataset and $m$ is the total number of model families. To prove identifiability, we adopt standard assumptions from the factor analysis literature (Chen et al., 2019) or regression literature, which assumes that the skills vectors $\theta^{(i)}\in{\mathbf{R}}^{1\times d}$’s are standardized, i.e., their average is null while their covariance matrix is fixed, $\text{rank}(\Lambda)=d$, and $\text{rank}(X)=p$.
Assumption B.1 (Identifiability constraints).
@@
One possible choice for the covariance matrix is $\Psi=I_{d}$ (Chen et al., 2019), which assumes uncorrelated skills. One implicit implication of Assumption B.1 is that $n\geq p\geq d$ must be satisfied, otherwise the covariance matrix cannot be full rank. This condition is satisfied in our experiments. Under Assumption B.1, we show the identifiability of the model parameters up to a transformation of $\Lambda$ tied to a transformation of $B$, which leaves the outputs of the model unchanged. This means that we can potentially approximate the true values for $\Lambda$ and $B$ up to a transformation, which is usually the norm within the class of exploratory factor analysis models.
Theorem B.2.
Given that the true set of model parameters is $(\Lambda,b,B)$, if there is another set of parameters $(\tilde{\Lambda},\tilde{b},\tilde{B})$ that satisfy
@@
then, under the Assumption B.1, we have $\tilde{b}=b$, $\tilde{\Lambda}=\Lambda M$, and $\tilde{B}=B(M^{\top})^{-1}$ for an invertible matrix $M\in{\mathbf{R}}^{d\times d}$. In particular, $M$ is orthogonal if $\Psi=I_{d}$, i.e., $M^{\top}M=MM^{\top}=I_{d}$.
@@
  • rank$(\Lambda^{\prime})=$rank$(\Lambda^{\prime}\Lambda^{\prime\top})=$rank$(\tilde{\Lambda}^{\prime}\tilde{\Lambda}^{\prime\top})=$rank$(\tilde{\Lambda}^{\prime})$. We see this by realizing that the null spaces of $\Lambda^{\prime\top}$ and $\Lambda^{\prime}\Lambda^{\prime\top}$ are the same: for an arbitrary vector $z$, $\Lambda^{\prime\top}z=0\Rightarrow\Lambda^{\prime}\Lambda^{\prime\top}z=0$ and $\Lambda^{\prime}\Lambda^{\prime\top}z=0\Rightarrow\Lambda^{\prime\top}\Lambda^{\prime}\Lambda^{\prime\top}z=0\Rightarrow\Lambda^{\prime\top}z=0$, where the last implication follows from the assumption that $\Lambda^{\prime\top}\Lambda^{\prime}$ is full rank ($\text{rank}(\Lambda^{\prime})=d$). Because the null spaces of $\Lambda^{\prime\top}$ and $\Lambda^{\prime}\Lambda^{\prime\top}$ are the same, their ranks should be the same as well. The same reasoning applies to $\tilde{\Lambda}^{\prime}\tilde{\Lambda}^{\prime\top}$ and $\tilde{\Lambda}^{\prime}$, proving this intermediate result.
@@
$\Lambda^{\prime}\Lambda^{\prime\top}=\tilde{\Lambda}^{\prime}\tilde{\Lambda}^{\prime\top}=\Lambda^{\prime}UU^{\top}\Lambda^{\prime\top}\Rightarrow\Lambda^{\prime\top}\Lambda^{\prime}\Lambda^{\prime\top}\Lambda^{\prime}=\Lambda^{\prime\top}\Lambda^{\prime}UU^{\top}\Lambda^{\prime\top}\Lambda^{\prime}\Rightarrow UU^{\top}=I$
@@
Appendix C Model fitting
$(\hat{\gamma},\hat{\sigma},\hat{b},\hat{\Lambda},\hat{\alpha},\hat{\beta})=\underset{\begin{subarray}{c}\gamma_{j}\in[0,1],\text{ for }j\in\mathcal{J}\\ \sigma_{j}:{\mathbf{R}}\to[0,1]\text{ increasing },\text{ for }j\in\mathcal{J}\\ b_{j}\in{\mathbf{R}},\text{ for }j\in\mathcal{J};\Lambda\in{\mathbf{R}}^{J\times d}\\ \alpha_{ik}\in{\mathbf{R}},\text{ for }i\in\mathcal{I}\text{ and }1\leq k\leq d\\ \beta_{k}\in{\mathbf{R}}^{3},\text{ for }1\leq k\leq d\end{subarray}}{\arg\min}~{}~{}\sum_{i\in\mathcal{I}}\sum_{(s,t)\in\mathcal{E}_{i}}\sum_{j\in\mathcal{J}}\ell_{\delta}(\mu_{ij}(s,t),Y_{ij})$
@@ Appendix C Model fitting
Appendix D Connections with factor analysis
Sloth is heavily inspired by (exploratory) factor analysis models. Factor analysis is a statistical technique used to identify underlying relationships between observed variables by reducing the data’s dimensionality (Bishop & Nasrabadi, 2006; Chen et al., 2019). It assumes that multiple observed variables are influenced by a smaller number of unobserved/latent variables called factors (skills $\theta^{(i)}$, in our case). These factors help explain the correlations among the observed variables. The method aims to model the observed variability and reveal the structure behind the data by estimating the factor loadings ($\Lambda$, in our case). The classical model assumes
@@ Appendix D Connections with factor analysis
where $Y_{i}$ is a vector of variables of interest and $\varepsilon_{i}$ is an error term. There are versions for the factor model in which a nonlinear model for $Y_{i}$ is assumed, e.g., in item response theory (IRT) (Reckase, 2006; Chen et al., 2019). It is usually the case that $\theta_{i}$ is estimated using a random effects model, i.e., practitioners place a prior distribution on $\theta_{i}$. In our work, we assume $\theta_{i}$ is given by a function of observable covariates and a family-specific intercept, which is fitted using data.
Appendix E Motivating the interaction term in Sloth
@@
Appendix F PCA approach formulation
We follow the ideas of Ruan et al. (2024) as closely as possible to create a prediction method. Moreover, we follow their code [^5] See https://github.com/ryoungj/ObsScaling. and apply PCA with the same set of hyperparameters. Assume we have a matrix of scores $Y\in[0,1]^{n\times J}$ in which columns represent benchmarks and each row represents a language model. We compute the covariance matrix of benchmark scores using $Y$ and then compute its eigenvector matrix $U$, where the columns give the ordered eigenvectors (from the highest eigenvalue to the lowest one). To reduce the dimensions of $Y$, we keep only the first $d$ columns of $YU$, resulting in $\tilde{Y}\in{\mathbf{R}}^{n\times d}$. For each column of $\tilde{Y}$ (principal components; PCs), we train a linear regression model using logFLOPs as the covariate; in this case, either the intercept or both the intercept and slope can be family-dependent. At test time, we predict the PCs of a held-out model and then go back to the original coordinate axis to obtain the final predictions by computing $\sum_{j=1}^{d}\hat{\text{PC}_{j}}U_{\cdot,j}\in{\mathbf{R}}^{J}$.
@@ Appendix H Extra performance prediction results
In this section, we present the full versions of the figures presented in the main text.
H.1 Test families have exactly one model in the training set
H.1.1 Average prediction loss across models
@@
H.1.2 Family-specific prediction losses
@@
H.2 Test families have exactly two models in the training set
H.2.1 Average prediction loss across models
@@
H.2.2 Family-specific prediction losses
@@ I.3 Results for $d=4$
In this section, we visually compare Sloth considering trainable and logistic link function $\sigma$, (Owen, 2024)’s model (“FLOPs (shared intercept)”) and our adaptation of Ruan et al. (2024)’s observational scaling law (“PCA + FLOPs”) described in Appendix F. For this experiment, we study the two Open LLM Leaderboards separately and consider LLaMa-3 and Yi-1.5 families as the test families; we make this choice because both families are popular ones and the training set size is the same for all models in each family, making comparison between models possible (in the x-axis, we use model size). For LLaMA-3, we just include one model from that family in the training set and do not train a family-specific slope for PCA+FLOPs. For Yi-1.5, we include two models in the training set and train a family-specific slope for PCA+FLOPs. In summary, we see that: (i) training the link function can produce a much more flexible scaling law that can better predict performance saturation (e.g., the performance of Yi-1.5 in ARC, HellaSwag etc.), (ii) training no family-specific parameters at all (“FLOPs (shared intercept)”) usually produce poor prediction results, and (iii) PCA+FLOPs often produces flatter curves that underestimate the performance of bigger models, e.g., see Yi-1.5 in TruthfulQA, GSM8k, and MMLU.
@@
Appendix L Comparing against Ruan et al. (2024) in their observational scaling law setting
In this section, we compare Sloth with Ruan et al. (2024)’s observational scaling law; that is, we extract abstract skills using a set of benchmark scores and then use those skills to predict the performance of models of interest in a target downstream task. For this experiment, we use the same data and tasks explored in Section 4.4. For our method, we fit Sloth using benchmark data from all models, including performance data of LLaMa-3-70B models, and extract the skills of each model. For Ruan et al. (2024)’s method, we fit PCA on the benchmark data to extract the skills. For both methods, we set $d=3$ and then fit a regression model with a logistic link to predict downstream performance from skills. Figures 36 and 37 present the prediction results for both methods and Figures 38 and 39 give the loading of both approaches. In both plots, out-sample prediction has a similar prediction error. At the same time, the in-sample fit is better for Sloth in the coding task and for Ruan et al. (2024)’s observational scaling law in the emotional intelligence task. Regarding the loading, it is possible to draw some similarities, e.g., the presence of instruction following skill, but there is no one-to-one correspondence between skills.
@@
Figure 36: Predicting code completion of LLaMa 3 70B (base/instruct) with Sloth vs Obs. Scaling Law (Ruan et al., 2024).
@@
Figure 37: Predicting emotional intelligence of LLaMa 3 70B (base/instruct) with Sloth vs Obs. Scaling Law (Ruan et al., 2024).

### v5

@@ 5 Conclusion
Limitations. From the predictive side of Sloth, we believe that the main limitation is that the model is still dependent, most of the time, on seeing data from at least one LLM from the LLM family of interest. Moreover, we train the link function in the best version of Sloth using flexible neural networks, which can interpolate data very well, but have no guarantee of extrapolation when the (hypothetical) LLM of interest is very different from others in the training set. From the interpretability side, we only understand the identification problems, such as transformations in the latent space, that can arise in a simple instance of Sloth: fixed activation function $\sigma$ and $\gamma_{j}$’s. This fact limits our understanding and interpretability of the most advanced versions of the model.
6 Acknowledgements
This paper is supported by the National Science Foundation (NSF) grants DMS-2027737, DMS-2113373, DMS-2414918, and a gift from OpenAI.
Appendix A Identifiability of model parameters and interpretability
To interpret Sloth parameters, we need to guarantee they are identifiable. Given that our scaling law models the function $\mu_{ij}(s,t)={\mathbf{E}}[Y_{ij}(s,t)]$, that condition is equivalent to the following statement: if two sets of parameters are responsible for characterizing $\mu_{ij}(s,t)$, then those set of parameters should be the same up to predictable variations such as translations or rotations. To prove identifiability, we work with a fixed and invertible $\sigma$, as usually done in the literature, and assume $\gamma_{j}$’s are fixed. The last condition is reasonable since these constants are usually known beforehand, e.g., it is well accepted that the lower asymptote $\gamma_{j}$ for MMLU [hendrycks2020Measuring] performance is $25\%$ which is given by $100\%$ divided by the number of multiple-choice alternatives. Denote our fixed design matrix as $X\in{\mathbf{R}}^{n\times p}$, where each row is given by an LLM and $p$ equals 3 plus the number of families, and define
@@ Appendix A Identifiability of model parameters and interpretability
such that the rows of $XB\in{\mathbf{R}}^{n\times d}$ give the skills vectors $\theta^{(i)}\triangleq(XB)^{(i)}$’s of all models in our dataset. Here $n$ denotes the total number of models in the dataset and $m$ is the total number of model families. To prove identifiability, we adopt standard assumptions from the factor analysis literature [chen2019joint] or regression literature, which assumes that the skills vectors $\theta^{(i)}\in{\mathbf{R}}^{1\times d}$’s are standardized, i.e., their average is null while their covariance matrix is fixed, $\text{rank}(\Lambda)=d$, and $\text{rank}(X)=p$.
Assumption A.1 (Identifiability constraints).
@@
One possible choice for the covariance matrix is $\Psi=I_{d}$ [chen2019joint], which assumes uncorrelated skills. One implicit implication of Assumption A.1 is that $n\geq p\geq d$ must be satisfied, otherwise the covariance matrix cannot be full rank. This condition is satisfied in our experiments. Under Assumption A.1, we show the identifiability of the model parameters up to a transformation of $\Lambda$ tied to a transformation of $B$, which leaves the outputs of the model unchanged. This means that we can potentially approximate the true values for $\Lambda$ and $B$ up to a transformation, which is usually the norm within the class of exploratory factor analysis models.
Theorem A.2.
[…]
@@
then, under the Assumption A.1, we have $\tilde{b}=b$, $\tilde{\Lambda}=\Lambda M$, and $\tilde{B}=B(M^{\top})^{-1}$ for an invertible matrix $M\in{\mathbf{R}}^{d\times d}$. In particular, $M$ is orthogonal if $\Psi=I_{d}$, i.e., $M^{\top}M=MM^{\top}=I_{d}$.
@@
[…]
@@
$\Lambda^{\prime}\Lambda^{\prime\top}=\tilde{\Lambda}^{\prime}\tilde{\Lambda}^{\prime\top}=\Lambda^{\prime}UU^{\top}\Lambda^{\prime\top}\Rightarrow\Lambda^{\prime\top}\Lambda^{\prime}\Lambda^{\prime\top}\Lambda^{\prime}=\Lambda^{\prime\top}\Lambda^{\prime}UU^{\top}\Lambda^{\prime\top}\Lambda^{\prime}\Rightarrow$
$\Rightarrow UU^{\top}=(\Lambda^{\prime\top}\Lambda^{\prime})^{-1}(\Lambda^{\prime\top}\Lambda^{\prime})(\Lambda^{\prime\top}\Lambda^{\prime})(\Lambda^{\prime\top}\Lambda^{\prime})^{-1}=I$
@@
A.1 Interpretability
@@
Appendix B Compute-optimal scaling
In this section, we derive compute-optimal scaling rules for skills. Specifically, given a language model family and a particular skill, our goal is to determine the model configuration that maximizes performance under a fixed computation budget $6st=c$.
Consider a model family $i$ with skill $k$ defined as
$\theta_{ik}(s,t)=\alpha_{ik}+\beta_{k0}\log(s)+\beta_{k1}\log(t)+\beta_{k2}\log(s)\log(t).$
Letting $u=\log(s)$, $v=\log(t)$, and $l=\log(c)-\log(6)$, we formulate the optimization problem as:
$\max_{u,v}\alpha_{ik}+\beta_{k0}u+\beta_{k1}v+\beta_{k2}uv\quad\text{s.t.}\quad u+v=l.$
This reduces to a simpler optimization problem in terms of $u$ alone:
$\max_{u}g_{ik}(u),\quad\text{where}\quad g_{ik}(u)\triangleq-\beta_{k2}u^{2}+(\beta_{k0}-\beta_{k1}+\beta_{k2}l)u+(\alpha_{ik}+\beta_{k1}l).$
To prevent our compute-optimal scaling method from making unreasonable predictions, we further restrict its solution to the ranges of $u$ and $v$ observed in the training data. Hence, we impose constraints $u\in[\underline{u},\overline{u}]$ and $v=l-u\in[\underline{v},\overline{v}]$, where bounds $\underline{u},\overline{u},\underline{v},\overline{v}$ are set based on quantiles from training data. Combining these constraints yields:
$u\in\mathcal{U}\triangleq[\max(l-\overline{v},\underline{u}),\min(l-\underline{v},\overline{u})].$
Thus, the optimization problem becomes:
$\max_{u\in\mathcal{U}}g_{ik}(u),$
which is straightforward to solve. Specifically, if $\beta_{k2}>0$, $g_{ik}(u)$ is a concave parabola, and the maximizer is either at the vertex (if it falls within $\mathcal{U}$) or at one of the interval endpoints $\max(l-\overline{v},\underline{u})$ or $\min(l-\underline{v},\overline{u})$.
Table 2 extends the results in Table 1 for all skills.
FLOPs (1e19) Reasoning Knowledge Instruction Following
Params (B) Tokens (T) Params (B) Tokens (T) Params (B) Tokens (T)
100 0.93 0.18 0.16 1.04 0.16 1.04
578 5.36 0.18 0.16 6.02 0.30 3.24
3346 30.98 0.18 0.37 15.0 0.72 7.78
19360 72.0 0.45 2.15 15.0 2.15 15.0
112005 72.0 2.59 12.44 15.0 12.44 15.0
648000 72.0 15.0 72.0 15.0 72.0 15.0
Table 2: Optimal allocation of parameters (B) and tokens (T) across skills for various compute budgets (FLOPs).
Appendix C Connections with factor analysis
Sloth is heavily inspired by (exploratory) factor analysis models. Factor analysis is a statistical technique used to identify underlying relationships between observed variables by reducing the data’s dimensionality [bishop2006pattern, chen2019joint]. It assumes that multiple observed variables are influenced by a smaller number of unobserved/latent variables called factors (skills $\theta^{(i)}$, in our case). These factors help explain the correlations among the observed variables. The method aims to model the observed variability and reveal the structure behind the data by estimating the factor loadings ($\Lambda$, in our case). The classical model assumes
@@ Appendix C Connections with factor analysis
where $Y_{i}$ is a vector of variables of interest and $\varepsilon_{i}$ is an error term. There are versions for the factor model in which a nonlinear model for $Y_{i}$ is assumed, e.g., in item response theory (IRT) [reckase200618, chen2019joint]. It is usually the case that $\theta_{i}$ is estimated using a random effects model, i.e., practitioners place a prior distribution on $\theta_{i}$. In our work, we assume $\theta_{i}$ is given by a function of observable covariates and a family-specific intercept, which is fitted using data.
Appendix D Motivating the interaction term in Sloth
@@
Appendix E PCA approach formulation
We follow the ideas of ruan2024observational as closely as possible to create a prediction method. Moreover, we follow their code [^6] See https://github.com/ryoungj/ObsScaling. and apply PCA with the same set of hyperparameters. Assume we have a matrix of scores $Y\in[0,1]^{n\times J}$ in which columns represent benchmarks and each row represents a language model. We compute the covariance matrix of benchmark scores using $Y$ and then compute its eigenvector matrix $U$, where the columns give the ordered eigenvectors (from the highest eigenvalue to the lowest one). To reduce the dimensions of $Y$, we keep only the first $d$ columns of $YU$, resulting in $\tilde{Y}\in{\mathbf{R}}^{n\times d}$. For each column of $\tilde{Y}$ (principal components; PCs), we train a linear regression model using logFLOPs as the covariate; in this case, either the intercept or both the intercept and slope can be family-dependent. At test time, we predict the PCs of a held-out model and then go back to the original coordinate axis to obtain the final predictions by computing $\sum_{j=1}^{d}\hat{\text{PC}_{j}}U_{\cdot,j}\in{\mathbf{R}}^{J}$.
Appendix F Sloth is parameter-efficient: parameter count analysis
When the number of model families $f$ is moderately large, Sloth actually uses fewer parameters than the top-performing baselines (both of which consider family-specific intercept or train the activation function). This is because of the assumed latent skill structure of Sloth. For example, with $d=3$ and we use $\mathcal{J}=12$ benchmarks:
  • Sloth: $12\cdot(d+2)+d\cdot(f+3)=69+3f$ parameters
  • “FLOPs” baseline: $12\cdot(f+3)=36+12f$ parameters
  • “Size and Tokens” baseline: $12\cdot(f+5)=50+12f$ parameters
So for $f\geq 4$, Sloth uses fewer parameters than either baseline.
@@ Appendix H Extra performance prediction results
In this section, we present the full versions of the figures presented in the main text and some other extra results.
H.1 Mean Absolute Percentage Error (MAPE) plot
Refer to caption Refer to caption
Figure 10: MAPE version of Figure 1. The results are given in percentage points, and we can see that the results are qualitatively similar to the MAE version, with Sloth producing the best predictions for both Open LLM Leaderboards.
H.2 Test families have exactly one model in the training set
H.2.1 Average prediction loss across models
@@
H.2.2 Family-specific prediction losses
@@
H.3 Test families have exactly two models in the training set
H.3.1 Average prediction loss across models
@@
H.3.2 Family-specific prediction losses
@@ I.3 Results for $d=4$
In this section, we visually compare Sloth considering trainable and logistic link function $\sigma$, owen2024predictable’s model (“FLOPs (shared intercept)”) and our adaptation of ruan2024observational’s observational scaling law (“PCA + FLOPs”) described in Appendix E. For this experiment, we study the two Open LLM Leaderboards separately and consider LLaMa-3 and Yi-1.5 families as the test families; we make this choice because both families are popular ones and the training set size is the same for all models in each family, making comparison between models possible (in the x-axis, we use model size). For LLaMA-3, we just include one model from that family in the training set and do not train a family-specific slope for PCA+FLOPs. For Yi-1.5, we include two models in the training set and train a family-specific slope for PCA+FLOPs. In summary, we see that: (i) training the link function can produce a much more flexible scaling law that can better predict performance saturation (e.g., the performance of Yi-1.5 in ARC, HellaSwag etc.), (ii) training no family-specific parameters at all (“FLOPs (shared intercept)”) usually produce poor prediction results, and (iii) PCA+FLOPs often produces flatter curves that underestimate the performance of bigger models, e.g., see Yi-1.5 in TruthfulQA, GSM8k, and MMLU.
@@
Appendix L Comparing against ruan2024observational in their observational scaling law setting
In this section, we compare Sloth with ruan2024observational’s observational scaling law; that is, we extract abstract skills using a set of benchmark scores and then use those skills to predict the performance of models of interest in a target downstream task. For this experiment, we use the same data and tasks explored in Section 4.4. For our method, we fit Sloth using benchmark data from all models, including performance data of LLaMa-3-70B models, and extract the skills of each model. For ruan2024observational’s method, we fit PCA on the benchmark data to extract the skills. For both methods, we set $d=3$ and then fit a regression model with a logistic link to predict downstream performance from skills. Figures 37 and 38 present the prediction results for both methods and Figures 39 and 40 give the loading of both approaches. In both plots, out-sample prediction has a similar prediction error. At the same time, the in-sample fit is better for Sloth in the coding task and for ruan2024observational’s observational scaling law in the emotional intelligence task. Regarding the loading, it is possible to draw some similarities, e.g., the presence of instruction following skill, but there is no one-to-one correspondence between skills.
@@
Figure 37: Predicting code completion of LLaMa 3 70B (base/instruct) with Sloth vs Obs. Scaling Law [ruan2024observational].
@@
Figure 38: Predicting emotional intelligence of LLaMa 3 70B (base/instruct) with Sloth vs Obs. Scaling Law [ruan2024observational].